# On the Low-Temperature MCMC threshold:

### v2

@@ 12.2.1. The Pure Noise Model
Lemma 12.12.
[New version of Lemma A.9 of [GZ22b]] For all $\rho\in(\rho_{*},1]$, $\zeta\in(0,1)$ and $t=\sqrt{1+\sigma^{2}/k}\exp\left(-(1-\alpha)\frac{k\log\frac{p}{k}}{n}\right)$, it holds that
$\mathbb{P}(\Upsilon_{\rho})\leq\left(\frac{2^{\frac{1}{2}}}{\zeta^{\frac{1}{12}}R_{1}^{\frac{1}{24}}}\right)^{n}$ (12.12)
Proof of Lemma 12.12.
@@
as desired, which proves Lemma 12.12. ∎
Lemma 12.13.
[New version of Lemma A.10 of [GZ22b]] Suppose that for some $\eta>0$, $k=o(p^{1/3-\eta})$. Then, for all $\rho\in[\frac{1}{k},\rho_{*}]$ and $\zeta\in(0,1)$, it holds that for large enough $k$,
@@
Proof of Lemma 12.13.
By Lemma 12.10, we have
@@
By Lemma 12.9 and because $\zeta<1$, $f$ is concave. Therefore the minimum value of $f$ on $[\frac{1}{k},\rho_{*}]$ is either $f(\frac{1}{k})$ or $f(\rho_{*})$, and thus
@@
We are now ready to finish the proof of Proposition 12.11. Combine the results of Lemma 12.12 and Lemma 12.13 to get that for all $\rho\in[\frac{1}{k},1]$ and $\zeta\in(0,1)$,
@@
It was shown earlier by the union bound (12.11) that
@@
Now we apply Lemma 12.12 and Lemma 12.13 to see that
$\mathbb{E}_{Y}\left(\min\{1,\Upsilon-1\}\right)$ $\leq\zeta^{n}+k\left[2^{\frac{n}{2}}\left(R_{1}^{\gamma}\zeta^{k}\right)^{-\frac{n}{12}}+2^{\frac{n}{2}}\left(R_{1}^{\frac{1}{2}}\zeta^{2}\right)^{-\frac{n}{12}}+2^{\frac{n}{2}}\left(R_{1}^{\frac{1}{24}}\zeta^{\frac{1}{12}}\right)^{-n}\right]$ (12.13)
To prove Proposition 12.11, it will therefore suffice to show that the sum in (12.13) is converging to 0 as $k\rightarrow\infty$. Let $\zeta=R_{1}^{-\frac{\gamma}{2k}}$ (which is less than 1 for large enough $D$), so that $\log\zeta=-\frac{\gamma}{2k}\log R_{1}$. Then:
@@
Next, we will bound the third term of (12.13), $2^{\frac{n}{2}}k\left(R_{1}^{\frac{1}{2}}\zeta^{2}\right)^{-\frac{n}{12}}$. Rewrite it as
@@
Finally, we want to bound the fourth term of (12.13), $k\left[2^{\frac{n}{2}}\left(R_{1}^{\frac{1}{24}}\zeta^{\frac{1}{12}}\right)^{-n}\right]$. This follows by an identical argument as for the third term of (12.13), which completes the proof of Proposition 12.11.
Now we can use Proposition 12.11 to prove Theorem 12.7.
Proof of Theorem 12.7.
@@
$\alpha k\log\frac{p}{k}\leq k\log\frac{p}{k^{2}}$ (12.14)
By assumption of Theorem 12.7, we have $n\leq\alpha k\log(\frac{p}{k})/(\log D_{0})$, and so
@@
$\left\{v\in\{0,1\}^{p}:\lVert v\rVert_{0}=k,\lVert Y-Xv\rVert_{\infty}\leq\sqrt{k}\sqrt{1+\sigma^{2}/k}\exp\left(-(1-\alpha)\frac{k\log\frac{p}{k}}{n}\right)\right\}$ (12.15)
@@
then our goal is to obtain a lower bound on $Z_{t_{0}\sqrt{k}}$, where $Z_{t_{0}\sqrt{k}}$ denotes the cardinality of (12.15). We let $D=D_{0}^{\frac{1}{2}}$, and then all the conditions of Proposition 12.11 are satisfied. We now require another improvement of a lemma from [GZ22b].
Lemma 12.14.
[New version of Lemma A.11 of [GZ22b]] The following bound holds with high probability with respect to $Y$ as $k$ increases:
@@
Recall from (12.4) that for any $t>0$ and for any $y$,
@@
as $k\rightarrow\infty$ by Proposition 12.11. This proves the claim that $Z_{t_{0}\sqrt{k},\infty}\geq\frac{1}{2}\mathbb{E}\left[Z_{t_{0}\sqrt{k},\infty}|Y\right]$ with high probability. We combine this result with Lemma 12.14 to conclude that w.h.p.,
@@
and therefore w.h.p., $Z_{t_{0}\sqrt{k},\infty}\geq\frac{1}{2}R^{\frac{n}{4}}\geq R^{\frac{n}{5}}$ for sufficiently large $D_{0}$, which concludes the proof of Theorem 12.7. ∎
12.2.2. The Linear model
@@
Then $X^{\prime}$ has i.i.d. $N(0,1)$ entries, and $Y^{\prime}$ has i.i.d. $N(0,(k-|S|)+\sigma^{2})$ entries, independently of $X^{\prime}$. This satisfies all the assumptions of Theorem 12.7 except that we need $k-|S|\leq k-|S|+\sigma^{2}\leq 3(k-|S|)$, or equivalently, $\sigma^{2}\leq 2(k-|S|)$. In general this will not hold, but in the special case of $S=\emptyset$ we can apply Theorem 12.7 to get this result:
Corollary 12.15.
[New version of Corollary B.1 of [GZ22b]] Suppose $\sigma^{2}\leq 2k$. For every sufficiently large constant $D_{0}$, if $4^{1+\eta}k^{3+\eta}\leq p$ for some $\eta>0$, and
@@
$n\leq\frac{\alpha k\log\left(\frac{p-k}{k}\right)}{\log 2D_{0}}$ (12.16)
@@
We can then apply Corollary 12.15 with $2D_{0}$ in place of $D_{0}$ to get that there are at least $R^{n/5}$ zero-overlap $v$’s that satisfy
$n^{-\frac{1}{2}}\lVert Y-Xv\rVert_{2}\leq R\sqrt{2k+\sigma^{2}}\exp\left(-\frac{k\log\left(\frac{p-k}{k}\right)}{n}\right)$ (12.17)
@@
$n^{-\frac{1}{2}}\lVert Y-Xv\rVert_{2}\leq R\sqrt{2k+\sigma^{2}}\exp\left(-\frac{(1-o(1))k\log\left(\frac{p-k}{k}\right)}{n}\right)$ (12.18)
@@
13. Acknowledgments

### v3

@@ A.2.1. The Pure Noise Model
Lemma A.13.
[New version of Lemma A.9 of [27]] For all $\rho\in(\rho_{*},1]$, $\zeta\in(0,1)$ and $t=\sqrt{1+\sigma^{2}/k}\exp\left(-(1-\alpha)\frac{k\log\frac{p}{k}}{n}\right)$, it holds that
$\mathbb{P}(\Upsilon_{\rho})\leq\left(\frac{2^{\frac{1}{2}}}{\zeta^{\frac{1}{12}}R_{1}^{\frac{1}{24}}}\right)^{n}$ (A.21)
Proof of Lemma A.13.
@@
as desired, which proves Lemma A.13. ∎
Lemma A.14.
[New version of Lemma A.10 of [27]] Suppose that for some $\eta>0$, $k=o(p^{1/3-\eta})$. Then, for all $\rho\in[\frac{1}{k},\rho_{*}]$ and $\zeta\in(0,1)$, it holds that for large enough $k$,
@@
Proof of Lemma A.14.
By Lemma A.11, we have
@@
By Lemma A.10 and because $\zeta<1$, $f$ is concave. Therefore the minimum value of $f$ on $[\frac{1}{k},\rho_{*}]$ is either $f(\frac{1}{k})$ or $f(\rho_{*})$, and thus
@@
We are now ready to finish the proof of Proposition A.12. Combine the results of Lemma A.13 and Lemma A.14 to get that for all $\rho\in[\frac{1}{k},1]$ and $\zeta\in(0,1)$,
@@
It was shown earlier by the union bound (A.20) that
@@
Now we apply Lemma A.13 and Lemma A.14 to see that
$\mathbb{E}_{Y}\left(\min\{1,\Upsilon-1\}\right)$ $\leq\zeta^{n}+k\left[2^{\frac{n}{2}}\left(R_{1}^{\gamma}\zeta^{k}\right)^{-\frac{n}{12}}+2^{\frac{n}{2}}\left(R_{1}^{\frac{1}{2}}\zeta^{2}\right)^{-\frac{n}{12}}+2^{\frac{n}{2}}\left(R_{1}^{\frac{1}{24}}\zeta^{\frac{1}{12}}\right)^{-n}\right]$ (A.22)
To prove Proposition A.12, it will therefore suffice to show that the sum in (A.22) is converging to 0 as $k\rightarrow\infty$. Let $\zeta=R_{1}^{-\frac{\gamma}{2k}}$ (which is less than 1 for large enough $D$), so that $\log\zeta=-\frac{\gamma}{2k}\log R_{1}$. Then:
@@
Next, we will bound the third term of (A.22), $2^{\frac{n}{2}}k\left(R_{1}^{\frac{1}{2}}\zeta^{2}\right)^{-\frac{n}{12}}$. Rewrite it as
@@
Finally, we want to bound the fourth term of (A.22), $k\left[2^{\frac{n}{2}}\left(R_{1}^{\frac{1}{24}}\zeta^{\frac{1}{12}}\right)^{-n}\right]$. This follows by an identical argument as for the third term of (A.22), which completes the proof of Proposition A.12.
Now we can use Proposition A.12 to prove Theorem A.8.
Proof of Theorem A.8.
@@
$\alpha k\log\frac{p}{k}\leq k\log\frac{p}{k^{2}}$ (A.23)
By assumption of Theorem A.8, we have $n\leq\alpha k\log(\frac{p}{k})/(\log D_{0})$, and so
@@
$\left\{v\in\{0,1\}^{p}:\lVert v\rVert_{0}=k,\lVert Y-Xv\rVert_{\infty}\leq\sqrt{k}\sqrt{1+\sigma^{2}/k}\exp\left(-(1-\alpha)\frac{k\log\frac{p}{k}}{n}\right)\right\}$ (A.24)
@@
then our goal is to obtain a lower bound on $Z_{t_{0}\sqrt{k}}$, where $Z_{t_{0}\sqrt{k}}$ denotes the cardinality of (A.24).
We let $D=D_{0}^{\frac{1}{2}}$, and then all the conditions of Proposition A.12 are satisfied. We now require another improvement of a lemma from [27].
Lemma A.15.
[New version of Lemma A.11 of [27]] The following bound holds with high probability with respect to $Y$ as $k$ increases:
@@
Recall from (A.13) that for any $t>0$ and for any $y$,
@@
as $k\rightarrow\infty$ by Proposition A.12. This proves the claim that $Z_{t_{0}\sqrt{k},\infty}\geq\frac{1}{2}\mathbb{E}\left[Z_{t_{0}\sqrt{k},\infty}|Y\right]$ with high probability. We combine this result with Lemma A.15 to conclude that w.h.p.,
@@
and therefore w.h.p., $Z_{t_{0}\sqrt{k},\infty}\geq\frac{1}{2}R^{\frac{n}{4}}\geq R^{\frac{n}{5}}$ for sufficiently large $D_{0}$, which concludes the proof of Theorem A.8. ∎
A.2.2. The Linear model
@@
Then $X^{\prime}$ has i.i.d. $N(0,1)$ entries, and $Y^{\prime}$ has i.i.d. $N(0,(k-|S|)+\sigma^{2})$ entries, independently of $X^{\prime}$. This satisfies all the assumptions of Theorem A.8 except that we need $k-|S|\leq k-|S|+\sigma^{2}\leq 3(k-|S|)$, or equivalently, $\sigma^{2}\leq 2(k-|S|)$. In general this will not hold, but in the special case of $S=\emptyset$ we can apply Theorem A.8 to get this result:
Corollary A.16.
[New version of Corollary B.1 of [27]] Suppose $\sigma^{2}\leq 2k$. For every sufficiently large constant $D_{0}$, if $4^{1+\eta}k^{3+\eta}\leq p$ for some $\eta>0$, and
@@
$n\leq\frac{\alpha k\log\left(\frac{p-k}{k}\right)}{\log 2D_{0}}$ (A.25)
@@
We can then apply Corollary A.16 with $2D_{0}$ in place of $D_{0}$ to get that there are at least $R^{n/5}$ zero-overlap $v$’s that satisfy
$n^{-\frac{1}{2}}\lVert Y-Xv\rVert_{2}\leq R\sqrt{2k+\sigma^{2}}\exp\left(-\frac{k\log\left(\frac{p-k}{k}\right)}{n}\right)$ (A.26)
@@
$n^{-\frac{1}{2}}\lVert Y-Xv\rVert_{2}\leq R\sqrt{2k+\sigma^{2}}\exp\left(-\frac{(1-o(1))k\log\left(\frac{p-k}{k}\right)}{n}\right)$ (A.27)
@@
A.3. Proof that structural conditions in [67] require $n=\Omega(k\sigma)$:
Here we show that the (deterministic) structural conditions in [67] are, with high probability, not satisfied for a design matrix $X$ with i.i.d. Gaussian entries when $n=o(k\sigma)$. Let $X$ be an $n\times p$ matrix with i.i.d. standard Gaussian entries. For any subset $\gamma\subseteq[p]$, denote by $X_{\gamma}$ the $n\times|\gamma|$ submatrix of $X$ given by the corresponding columns of $X$. For any $\gamma$, let $\Phi_{\gamma}$ be the orthogonal projection operator onto the span of the column space of $X_{\gamma}$. Let $Z\sim N(0,I_{n})$ independent of $X$.
Lemma A.17.
Let $n=\omega(k)$. It holds with high probability as $k\rightarrow\infty$ that
$\mathbb{E}\left[\max_{|\gamma|=k}\max_{j\in[p]\setminus\gamma}\frac{1}{\sqrt{n}}\left|\left\langle(I-\Phi_{\gamma})X_{j},Z\right\rangle\right|\right]=\Omega\left(\sqrt{\frac{k^{2}}{n}}\right)$ (A.28)
Consequently, Assumption B in [67] (the Sparse Projection Condition) requires $L$ to be at least $\Omega(k^{2}/(n\log p))$ to hold with high probability under our setting. The High SNR condition (equation (9a) in that work) and the fact that we take $v^{*}\in\{0,1\}^{p}$ implies that their mixing time result (i.e. Theorem 2 of [67]) requires $n=\Omega(L\sigma^{2}\log p)$, or equivalently, $n=\Omega(k\sigma)$.
Proof of Lemma A.17.
Define the random set $T\subseteq\mathbb{R}^{n}$ as
$T:=\left\{\frac{1}{\sqrt{n}}\left(I-\Phi_{\gamma}\right)X_{j}:\gamma\subseteq[p],|\gamma|=k,j\in[p]\setminus\gamma\right\}$
By Sudakov’s minorization inequality (see e.g., Corollary 7.4.3 of [63]), it holds for any $\varepsilon>0$ that
$\mathbb{E}\left[\max_{t\in T}\left\langle t,Z\right\rangle\right]=\Omega\left(\varepsilon\sqrt{\log\mathcal{N}(T,\varepsilon)}\right)$ (A.29)
where $\mathcal{N}(T,\varepsilon)$ is the $\varepsilon$-covering number of $T$. It therefore suffices to show that with high probability, at least $e^{\Omega(k)}$ points in $T$ are pairwise separated by at least $\sqrt{\frac{k}{n}}$, as then each of these points will require its own $\sqrt{\frac{k}{n}}$-sphere when forming a covering.
Fix a choice of $j$. Fix any $\gamma\neq\gamma^{\prime}$ satisfying $|\gamma\cap\gamma^{\prime}|\geq k/3$ and $j\notin\gamma\cup\gamma^{\prime}$. We first note that since $X_{j}$ is standard Gaussian independent of $X_{\gamma}$ and $X_{\gamma^{\prime}}$, it holds that
$\mathbb{E}\left(\left\lVert\frac{1}{\sqrt{n}}\left(I-\Phi_{\gamma}\right)X_{j}-\frac{1}{\sqrt{n}}\left(I-\Phi_{\gamma^{\prime}}\right)X_{j}\right\rVert^{2}\right)=\mathbb{E}\left(\left\lVert\frac{1}{\sqrt{n}}\left(\Phi_{\gamma^{\prime}}-\Phi_{\gamma}\right)X_{j}\right\rVert^{2}\right)$
$=$ $\mathbb{E}\left(\mathbb{E}\left(\left\lVert\frac{1}{\sqrt{n}}\left(\Phi_{\gamma^{\prime}}-\Phi_{\gamma}\right)X_{j}\right\rVert^{2}\right)|X_{\gamma},X_{\gamma^{\prime}}\right)$
$=$ $\mathbb{E}\left(\frac{1}{n}\left\lVert\Phi_{\gamma^{\prime}}-\Phi_{\gamma}\right\rVert_{F}^{2}\right)$
$=$ $\mathbb{E}\left(\frac{1}{n}\left(\left\lVert\Phi_{\gamma^{\prime}}\right\rVert_{F}^{2}+\left\lVert\Phi_{\gamma}\right\rVert_{F}^{2}-2\mathrm{tr}\left(\Phi_{\gamma^{\prime}}\Phi_{\gamma}\right)\right)\right)$
$=$ $\frac{1}{n}\left(k+k-2|\gamma\cap\gamma^{\prime}|\right)=2\frac{k-|\gamma\cap\gamma^{\prime}|}{n}$
where the last line holds because $\Phi_{\gamma}$ and $\Phi_{\gamma^{\prime}}$ are orthogonal projection matrices.
Next, let $C>0$ be a large constant. Then, with probability at least $1-e^{-n}$, it holds that $\left\lVert X_{j}\right\rVert\leq C\sqrt{n}$. Condition on this event. We next observe that the function $x\rightarrow\left\lVert\frac{1}{\sqrt{n}}(\Phi_{\gamma^{\prime}}-\Phi_{\gamma})x\right\rVert^{2}$ is $\frac{2C}{\sqrt{n}}$-Lipschitz when restricted to the ball of radius $C\sqrt{n}$. Therefore, by Gaussian concentration of Lipschitz functions, it holds for some universal constant $c>0$ and any $t>0$ that
$\mathbb{P}\left(\left|\left\lVert\frac{1}{\sqrt{n}}(\Phi_{\gamma^{\prime}}-\Phi_{\gamma})X_{j}\right\rVert^{2}-2\frac{k-|\gamma\cap\gamma^{\prime}|}{n}\right|\geq t\right)\leq 2\exp^{-cnt^{2}}$
Since we have $|\gamma\cap\gamma^{\prime}|\leq k/3$ assumption, if we then choose $t=\sqrt{k/n}$, then we see that with probability at least $1-2\exp\left(-ck\right)$, it holds that
$\left\lVert\frac{1}{\sqrt{n}}(\Phi_{\gamma^{\prime}}-\Phi_{\gamma})X_{j}\right\rVert^{2}\geq\frac{k}{n}$
By a union bound over $\gamma$ and $\gamma^{\prime}$, this holds with high probability for at least $\exp(c^{\prime}k)$ choices of $(\gamma,\gamma^{\prime})$ satisfying $|\gamma\cap\gamma^{\prime}|\leq k/3$, where $c^{\prime}<c$ is a small enough constant (such a set of subsets of $[p]$ is guaranteed to exist by Lemma 4.14 of [56]). Therefore, with high probability we have that the $\sqrt{\frac{k}{n}}$-covering number of $T$, $\mathcal{N}\left(T,\sqrt{\frac{k}{n}}\right)$, is at least $e^{\Omega(k)}$. The desired result then follows by (A.29). ∎